OpenStudy (anonymous):

log sub 8 (√x^5/y^3z^6)

OpenStudy (zehanz):

So:\[\log_{8}\frac{ \sqrt{x^5} }{ y^3z^6 } \]Use the following log-rules:\[\log_{n}\frac{ a }{ b } =\log_{n}a-\log_{n}b\]\[\log_{n}a^k=k \cdot \log_{n}a \]